Skip to end of metadata
Go to start of metadata

Sonics and gas analyzers

tse07 10m sonic

Replaced and working.

rsw02 20m

Still down, still needs to be replaced.

v07 4m

Still suspicious. A plot of ldiag.4m.v07 over the last several days is below, and the bad periods look diurnal. Maybe a temperature problem? This is an RMYoung, so I can look for a working RMYoung spare with which to replace it.

v06 20m LiCOR

Still missing samples later in the day, but it doesn't look that bad judging from a plot of the last several days. I think lidiag is usually non-zero, so the only diagnostic that matters is how many of co2.20m.v06 and h2o.20m.v06 are missing.  It does look like it's getting worse, but I don't know of anything we can do about it since we have no spares.

TRH Sensors

tse04 and tse11 60m

INEGI will replace at least tse04 tomorrow.

Other issues

tnw08 Vdsm

Still not there.  Steve says we have no more voltage motes, so it will never be there.

LiCORs for rsw01 and rsw07

The ARL shipment did not arrive today. Chris says ARL is going to plan as if they won't arrive and make do with the instruments they have.

 

  • No labels

2 Comments

  1. Gary Granger AUTHOR

    I have a replacement for the RMYoung at 20m on rsw02 tested and ready to go, using the one that was originally removed from rsw02 20m.  Orson will help us replace it tomorrow.

     

  2. Gary Granger AUTHOR

    For v07 4m, there is only one more RMYoung left which is not marked bad.  I've tested it, and in the ops center the diagnostic bits are set maybe one out of every 30 samples.  Whereas the sonic already on v07 4m is solid good at the moment, it's only during the day that it looks like it goes almost completely bad.

    So as far as fixing the sonic already at the site, I could jiggle some wires and check for loose connections in the terminal block, but the only one that I suppose could be related to these symptoms is power.  A plot of ldiag versus Vdsm does suggest that power could be related.  When power is above 12.6V or so, sonic is good.  What's happening that power is up at night and down during the day?  Is there something else at that site drawing power during the day?  Could it be the adapter performance declines with heat?  Maybe this site doesn't have a 14V adapter but should?  Do we have any more 14V adapters?  Maybe it's just a flaky adapter that didn't completely fail in a lightning storm but still needs to be swapped out.  The LiCOR at v07 has also been problematic for a long time, maybe it's having problems related to the power also.